# A First Look at Package-to-Group Mechanism: An Empirical Study of the Linux Distributions

### v1

@@ V-A Methodology
Targeting the above goal, our method consists of two parts: a summary of group change patterns and an analysis of package flow directions.
@@ V-A1 Change Patterns
We manually review the open-source repositories of the Linux distributions mentioned in Section III-B to locate the URLs of the source files containing group and package information. Then we use the request library to retrieve group information for each release version. Finally, we utilize the group information to calculate the added and removed groups between consecutive versions, including their names, functional descriptions, and contained packages.
To summarize these change patterns, we employed a manual annotation process conducted by a team comprising one PhD candidate and two master’s students majoring in computer science. All of them are fluent in English and have completed a semester course on operating systems. To guarantee the quality of the annotation process, we set a two-round labeling process. In the first round, each annotator independently reviewed the added and removed groups, labeling each change with a descriptive pattern. Based on these individual annotations, the team developed a taxonomy of change patterns. In the second round, the taxonomy is relabeled by two master students. In cases of disagreement, the PhD candidate was consulted to reach an agreement.
@@ V-A2 Flow Analysis
Packages in Linux distributions can be divided into two categories: those using the P2G and those that do not. Thus, the flow sources of packages adopting the P2G consist of two types: packages that previously did not use P2G but now do (i.e., S1) and new packages added to the Linux distribution (i.e., S2). The flow outcomes also fall into two categories: packages that remain in Linux but no longer use the P2G (i.e., O1) and packages that are completely removed from Linux (i.e., O2).
To calculate the number of packages in each scenario, we apply the following algorithm. For each Linux distribution version in Section III-B, we first identify the lists of added and removed packages (i.e., $ap,rp$) under P2G compared to the previous version, as well as the total number of packages in both the previous and current versions. For each package in the $ap$, we check whether it exists in the total packages of the previous version. if it does, it belongs to S1, and if not, it falls under S2. For each package in $rp$, we verify its presence in the total packages of the current version. if it is found, it belongs to O1, otherwise, it is classified as O2. Using this algorithm, we can calculate the number of packages in four scenarios for each version change. Then we merge them to obtain the final package flow outcomes.
@@ V-B1 Change Patterns
Figure 6 presents the taxonomy of change patterns in Linux distributions. There are three primary patterns for adding groups. ❶ Split Groups: This typically occurs when a group is too large or its functions too concentrated, making management and updates inconvenient. For instance, the Chinese Support group was split into two distinct groups(i.e., Simplified Chinese and Traditional Chinese) from Fedora 14 to Fedora 15. ❷ Add Features: As technology advances and new requirements arise, additional groups are created to accommodate these changes. This pattern includes 12 subtypes, including language support, font support, file system management, etc. For example, Fedora increased support for different languages and regions to enhance the internationalization from version 4 to 5. ❸ Rename Groups: This generally reflects technological changes or a shift in marketing strategy to ensure the group name better aligns with its function. For example, Fedora renamed OpenOffice.org Development group to LibreOffice Development. In addition, there are three patterns for removing groups. ❹ Remove Features: This can be categorized into two subtypes. The first occurs when a specific technology becomes outdated or is replaced by more efficient solutions, such as removing MeeGo NerBook UX Environment. The second subtype involves the elimination of features that are no longer necessary, such as support for a specific language or font. ❺ Merge Groups: To optimize resources and streamline management, groups with related or overlapping functionalities may be merged. For example, the Virtualization Client group and Virtualization Hypervisor group were combined into a unified Virtualization group. ❻ Replace Features: a group may be replaced to cover a broader or more updated technological scope. For instance, the Moblin Desktop Environment group is integrated into the more comprehensive Virtualization group.
To ensure the final taxonomy is comprehensive and representative, we conducted a validation survey involving 12 Linux practitioners and 2 researchers. We list each evolution pattern with its description and examples, inviting participants to indicate whether they agree with it. The results of the survey are summarized in Table II. For each pattern, we report the percentage of “yes” and “no” answers to the question asking whether participants agreed on it. The most approved pattern is merging existing groups, with 100% of “yes” answers. The least approved category is rename groups, which has been agreed by 40% of the survey participants (a non-negligible fraction of all the participants). Overall, the average rate of “yes” answers across all patterns is 70%, indicating that the final patterns align with the experience of most participants (only one pattern is below 50%).
@@
Remove Group Remove Features 60% 40%
@@ V-B2 Flow Analysis
Figure 7 shows the percentage of each flow situation. ❶ We can see that The number of packages added to groups surpasses those removed. Added packages from both flow sources make up 51.6%, while removed packages account for 48.4%. This suggests that with each version update, more packages are starting to adopt the P2G. ❷ Two types of flow sources for the P2G have the same percentage. Half of the packages come from the Linux distributions themselves (i.e., S1) and the other half comes from external sources (i.e., S2). ❸ Packages that no longer use the P2G are more likely to remain in Linux distributions. Among these, 29.4% have stayed in Linux distributions (i.e., O1) while 19% have been directly removed from them (i.e., O2).
Summary for RQ2: The P2G in Linux distributions has six major change patterns, including splitting and merging groups, etc. With version evolution, more packages adopt P2G, and even those stop using it often remain in Linux distributions.
@@ VI RQ3:Quality Assessment
We aim to evaluate the quality of groups under the P2G mechanism in the latest Linux distributions and identify potential issues.
@@ VI-A1 Proposed Metrics
  • Relevance: The description of a group is critical in guiding developers to choose packages, so it must accurately summarize the functionality of the packages.
@@
  • Distributions: the number of packages in a group should be neither too large nor too small. Overcrowded groups can lead to confusion in the discovery of specific items, while groups with two few packages (e.g., just one) may result in inefficient management.
Based on these expert criteria outlined above, we propose GValue, a metric for evaluating the quality of groups in OSS projects. Figure 8 provides an overview of our GValue, which consists of three steps. In the first step, two types of information are extracted from the Linux distributions repositories, which is package attributes and group attributes. In the second step, these extracted attributes are processed and transformed into a set of values that reflect the quality of groups from four perspectives, which are compact, relevance, difference, and distribution. Finally, these values are consolidated into a single metric. The following sections provide a detailed description of each step.
Extract raw data. As illustrated in figure 8, the step aims to extract group and package attributes from the Linux distribution repositories. The step first extract $n$ groups from Linux repositories, denoted as $O=\{g_{1},g_{2},...,g_{n}\}$. Then, three attributes are extracted for each group $g_{i}$, including group name, group description, and package list, expressed as $g_{i}=\{gn_{i},gd_{i},pl_{i}\}$. Assuming group $g_{i}$ contains $m$ packages, the package list is represented as $pl_{i}=\{p_{i1},p_{i2},...,p_{im}\}$. Additionally, we extract package attributes for each package in groups, including package name and package description, which can be defined as $p_{ij}=\{pn_{ij},pd_{ij}\}$.
Compactness Value. Compact refers to the functional similarity or dependency degree among packages within a group. To measure similarity, we leverage the rich semantic information contained in package function descriptions. Specifically, we utilize a pre-trained language model ($PLM$) to convert these descriptions into semantic vectors and calculate cosine similarity between them as similarity scores among packages.
@@
To quantify the dependency degree, we measure it by the path length in the package dependency graph. This approach effectively reflects the relationships between packages, as a shorter distance between nodes in the graph indicates a higher degree of dependency.
@@
Relevance Value. To calculate the relevance between a group’s description and the descriptions of packages in the group, we first input the group description and package descriptions into a pre-trained language model to obtain their semantic representation (i.e., $V_{i}=PLM(gd_{i})$). Then, we compute the similarity between the group description and the description of each package in this group (i.e., equation 1). Last, we take the average of these similarity scores as the overall relevance value for a group.
@@
Differentiation Value. A group contains three main attributes: name, description, and a list of packages. To comprehensively assess the differences between groups, we evaluate each attribute individually before combining them into an overall score. Specifically, we design methods to quantify the differences for each attribute individually, which are then integrated to provide a comprehensive assessment.
Firstly, edit distance is chosen to measure the difference between group names. This choice stems from our observation that some group names are simply character strings, making it difficult to extract meaningful semantic information, such as the group KDE in CentOS 7.
@@
where $ED(gn_{i},gn_{j})$ and $ndif(g_{i},g_{j})$ represent the edit distance and difference score between the name of i-th group and the j-th group, respectively. $\|gn_{i}\|$ denotes the string length of the i-th group’s name and $ndif(g_{i})$ refers to the different score between the name of i-th group and all other groups.
To assess the differences between group descriptions, we also use a pre-trained model to compute the semantic vector ($V_{i}$) for each group and measure the difference using cosine similarity. Since cosine similarity ranges from -1 to 1, where larger values indicate smaller differences, we apply interval shifting to normalize the values, restricting them to a range between 0 and 1 [34].
@@
Lastly, since the package list in a group can be viewed as a set, set similarity can be used as a metric to measure the difference between package lists. However, considering that the packages have varying levels of necessity(e.g., mandatory, default, and optional), treating all packages as equally important and applying set similarity directly is not appropriate. To address this, different weights are assigned to each type of package. The principle behind the weight assignment is that more essential a package is, the higher its weight. $w_{i}(p_{j})$ is the weight function for the packages. The determination of this function is based on multiple iterations through computing its correlation with manual scoring.
@@
The Jaccard similarity [35] is a statistical measure used to compare the similarity and diversity between finite sets. Building on this, we propose a quantitative method that incorporates both the Jaccard similarity and the above weight function.
@@
where $pdif(pl_{i},pl_{j})$ represents the difference score between the packages list in the i-th and j-th group. $pdiff(g_{i})$ denotes the difference between the package list in the i-th group and all other groups.
@@
GValue Fuse. Based on the calculations from the four dimensions mentioned above, we use the average value (i.e., $GValue(g_{i})$) as the comprehensive evaluation metric for assessing the quality of a group.
@@ VI-B1 Metrics Validation
We validated our GValue metric by evaluating five groups randomly selected from five different Linux distributions. The ground truth for these groups was established through a survey completed by 16 skilled participants. Then we calculate the Spearman correlation [37] between the human score and GValue score as an evaluation metric. The result is shown in Figure 9. As can be seen, the GValue have a strong correlation with human score ($rs>0.7$) in each aspect.
@@ VI-B2 Problems Analysis
Summary for RQ3: We developed a multi-dimensional feature-based method to evaluate group quality, which shows a strong correlation with human assessments. Our analysis of Linux distributions revealed that 16% of the groups exhibit poor quality, with two types of issues identified in the adoption of P2G.
@@ VII-A Methodology
VII-A1 Topic Analysis on group descriptions.
Latent Dirichlet Allocation [38] (LDA) is a widely used unsupervised learning method for extracting topics from large-scale unstructured text. By mining the latent semantic structure of the text, LDA enables topic modeling across a collection of descriptions to identify common themes. We apply LDA to conduct topic analysis on the group descriptions. Specifically, we start with the preprocessing of the group descriptions, including tokenizing and removing stopwords using NLTK [39]. We then construct a vocabulary dictionary using the Gensim [40] and generate a group-word frequency matrix, representing the frequency of each word in each group. Finally, each topic is characterized by the top ten words with the highest weights, revealing the latent semantic structure in the group descriptions. Additionally, as LDA requires specifying the number of topics in advance, we perform a series of experiments with various numbers of topics and evaluate them using the coherence score to determine the optimal number.
@@ VII-A2 Extraction of Key Information from package descriptions
We extract and compare the key information from the descriptions of the packages in groups with those that are not. Package descriptions often contain many redundant and trivial words, so we first perform tokenization and stopword removal. We then use the term frequency-inverse document frequency (TF-IDF) algorithm [41] to mine keywords. The TF-IDF algorithm determines the importance of a word by comparing its frequency in a specific package description against its rarity across all package descriptions. Essentially, the algorithm measures how relevant a word is within a given package description. Specifically, for a set of package descriptions $S$, a specific package description $p$ from this set, and a word $w$ in $p$, the relevance $i_{w}$ of the word is calculated as follows:
@@
where $f_{w,p}$ represents the frequency of $w$ in $p$, $\|S\|$ is the total number of package descriptions in the set $S$, and $f_{w,S}$ denotes the frequency of $w$ across the entire set $S$. Using this algorithm, we extract the most relevant words from each package description. To help Linux developers determine whether to apply the P2G mechanism to a package, we collect and compare the keywords from package descriptions both within and outside of groups.
@@ VII-B1 Topic in group descriptions
we set the number of topics between 1 and 20 and calculate their coherence scores. We find that the coherence score is the highest in the 3 topics, achieving a value of 0.48. Figure 10 presents the results for group descriptions under this situation. We can observe that the group descriptions mainly involve the following broad categories: Applications (e.g., fonts and tools), Development (e.g., programming and libraries), Graphical Desktop (e.g., desktop and KDE), Networks (e.g., network and virtualization), and Servers (e.g., servers). This suggests that developers tend to adopt P2G when a package’s functionality aligns with one of these categories.

### v2

@@ V-A Methodology
Targeting the above goal, our method consists of two parts: a content-based summary of group change patterns and a quantitative analysis of package flow directions.
@@ V-A1 Change Patterns
We manually review the open-source repositories of the Linux distributions mentioned in Section III-B to locate the URLs of the source files containing group and package information. Then we use the request library to scrape group information for each release version. Finally, we utilize the group information to calculate the added and removed groups between consecutive versions, including their names, functional descriptions, and contained packages.
To summarize the change patterns, we employ a manual annotation process conducted by a team. The team consists of one PhD candidate and two master students majoring in computer science. All of them are fluent in English and have completed a semester course on operating systems. To guarantee the quality of the annotation process, we set a two-round labeling process. In the first round, annotators independently review the added and removed groups, labeling each change with a descriptive pattern. Based on these individual annotations, the team can develop a taxonomy of change patterns. In the second round, the taxonomy is relabeled by two master students. In cases of disagreement, the PhD candidate is consulted to reach an agreement.
@@ V-A2 Flow Analysis
Packages in Linux distributions can be divided into two categories: those using P2G and those that do not. Thus, the flow sources consist of two types: packages that previously did not use P2G but now do (i.e., S1) and new packages added to the distributions (i.e., S2). The flow outcomes also fall into two categories: packages that remain in Linux but no longer use P2G (i.e., O1) and packages that are completely removed from distributions (i.e., O2).
To calculate the number of packages in each scenario, we apply the following algorithm. For each distribution version, we identify the added and removed packages (i.e., $ap,rp$) under P2G compared to its previous version, as well as the total number of packages in the previous and current versions. For each package in $ap$, we check whether it exists in the previous version. if it does, it belongs to S1, and if not, it falls under S2. For each package in $rp$, we verify its presence in the current version. If it is found, it belongs to O1, otherwise, it is classified as O2. Using this algorithm, we can calculate the number of packages in four scenarios for each version. Finally, we merge them to obtain the package flow results.
@@ V-B1 Change Patterns
Figure 6 presents the taxonomy of change patterns in Linux distributions. There are three primary patterns for adding groups. ❶ Split Groups: This typically occurs when a group is too large or its functions too concentrated, making management inconvenient. For instance, the Chinese Support group is split into two distinct groups(i.e., Simplified Chinese and Traditional Chinese) from Fedora 14 to Fedora 15. ❷ Add Features: As technology advances and new requirements arise, additional groups are created to accommodate these changes. This pattern includes 12 subtypes, e.g., language support, font support, and file system management. For example, Fedora increases support for different languages and regions to enhance internationalization from version 4 to 5. ❸ Rename Groups: This generally reflects technological changes or a shift in marketing strategy to ensure the group name aligns with its new function. For example, Fedora renamed OpenOffice.org Development group to LibreOffice Development. In addition, there are three patterns for removing groups. ❹ Remove Features: This can be categorized into two subtypes. The first occurs when a specific technology becomes outdated or is replaced by more efficient solutions, such as removing MeeGo NerBook UX Environment. The second subtype involves the elimination of features that are no longer necessary, such as support for a specific language or font. ❺ Merge Groups: To optimize resource management, groups with related or overlapping functionalities can be merged. For example, the Virtualization Client group and Virtualization Hypervisor group are combined into a unified Virtualization group. ❻ Replace Features: a group may be replaced to cover a broader or more updated technological scope. For instance, the Moblin Desktop Environment group is integrated into the more comprehensive Virtualization group.
To ensure the taxonomy is comprehensive and representative, we conducted a validation survey involving 12 Linux practitioners and 2 researchers. We list each pattern with its description and two examples, inviting participants to answer whether they agree with it. The results are summarized in Table II. For each pattern, we report the percentage of “yes” and “no” answers. The most approved pattern is merging existing groups, with 100% of “yes” answers. The least approved category is rename groups, which has been agreed by 40% of these participants (a non-negligible fraction of all the participants). Overall, the average rate of “yes” answers across all patterns is 70%, indicating that the patterns align with the experience of most participants (only one pattern is below 50%).
@@
Remove Groups Remove Features 60% 40%
@@ V-B2 Flow Analysis
Figure 7 shows the percentage of each flow situation. ❶ We can see that the number of packages added to groups surpasses those removed. Added packages from both flow sources make up 51.6%. This suggests that with each version update, more packages are starting to adopt P2G. ❷ Two types of flow sources for P2G have the same percentage. Half of the packages come from the Linux distributions themselves (i.e., S1) and the other half comes from external sources (i.e., S2). ❸ Packages that no longer use P2G are more likely to remain in distributions. Among these, 29.4% have stayed in distributions (i.e., O1) while 19% have been directly removed from them (i.e., O2).
Summary for RQ2: The P2G in Linux distributions has six major change patterns, including splitting and merging groups, etc. With version evolution, more packages adopt P2G and those no longer using P2G often remain in distributions.
@@ VI RQ3:Quality Assessment
We aim to evaluate the quality of groups under the P2G mechanism in the latest distributions and identify potential issues.
@@ VI-A1 Proposed Metrics
  • Relevance: The description of a group is critical in guiding developers to choose packages, so it should summarize the functionality of the packages accurately.
@@
  • Distributions: the number of packages in a group should be balanced. Overcrowded groups can lead to confusion when retrieving specific items, while groups with too few packages may result in inefficient management.
Based on these expert criteria, we propose GValue, a metric for evaluating the quality of groups in P2G. Figure 8 provides an overview of our GValue, which consists of three steps. In the first step, two types of information are extracted from the distribution repositories, which is package attributes and group attributes. In the second step, they are processed and transformed into a set of values that reflect the quality of groups according to the four expert criteria. Finally, these values are consolidated into a single metric. The following sections provide a detailed description for each step.
Extract raw data. As illustrated in figure 8, the step aims to extract group and package attributes from the distribution repositories. We first scrape $n$ groups from repositories, denoted as $O=\{g_{1},g_{2},...,g_{n}\}$. Then, three attributes are extracted for each group $g_{i}$, including name, description, and package list, expressed as $g_{i}=\{gn_{i},gd_{i},pl_{i}\}$. Assuming group $g_{i}$ contains $m$ packages, the package list is represented as $pl_{i}=\{p_{i1},p_{i2},...,p_{im}\}$. Additionally, we extract package attributes for each package in groups, including package name and package description, which can be defined as $p_{ij}=\{pn_{ij},pd_{ij}\}$.
Compactness Value. Compact refers to the functional similarity or dependency degree among packages within a group. To measure similarity, we leverage the rich semantic information contained in package function descriptions. Specifically, we utilize a pre-trained language model ($PLM$) to convert these descriptions into semantic vectors and calculate cosine similarity between them as similarity scores.
@@
Then, we quantify the dependency degree by the path length in the package dependency graph. It can reflect the relationships between packages effectively, as a shorter path in the graph indicates a higher degree of dependency.
@@
Relevance Value. To calculate the relevance between the descriptions of a group and packages in the group, we first input the group description and package descriptions into $PLM$ to obtain their semantic representation (i.e., $V_{i}=PLM(gd_{i})$). Then, we compute the cosine similarity between them (i.e., equation 1). Last, we take the average of these similarity scores as the overall relevance value for a group.
@@
Differentiation Value. A group has three main attributes: name, description, and a list of packages. To comprehensively assess the differences between groups, we evaluate each attribute individually before combining them into an overall score. Specifically, we design methods to quantify the differences for each attribute individually, which are then integrated to provide a comprehensive assessment.
Firstly, edit distance is chosen to measure the difference between names. This choice stems from our observation that some group names are simply character strings, making it difficult to extract semantic information, such as the group KDE in CentOS 7.
@@
where $ED(gn_{i},gn_{j})$ and $ndif(g_{i},g_{j})$ represent the edit distance and difference score between the i-th group’s name and the j-th group’s name, respectively. $\|gn_{i}\|$ denotes the string length of the i-th group’s name and $ndif(g_{i})$ refers to the different score between the name of i-th group and all other groups.
To assess the differences between group descriptions, we also use $PLM$ to compute the semantic vector ($V_{i}$) for each group and measure the difference using cosine similarity. Since cosine similarity ranges from -1 to 1 and larger values indicate smaller differences, we apply interval shifting to normalize the values, restricting them to a range between 0 and 1 [34].
@@
Lastly, since the package list in a group can be viewed as a set, set similarity can be used as a metric to measure the difference between package lists. However, considering that the packages have varying levels of necessity(e.g., mandatory, default, and optional), treating all packages as equally important and applying set similarity directly is not appropriate. To address this, different weights are assigned to each type of package. The principle behind the weight assignment is that the more essential a package is, the higher its weight. $w_{i}(p_{j})$ is the weight function for the packages. The determination of this function is based on multiple iterations through computing its correlation with manual scoring.
@@
Based on the weight function and Jaccard similarity [35], we propose a quantitative method to measure the differences among package lists in distinct groups.
@@
where $pdif(pl_{i},pl_{j})$ represents the difference score between the package list in the i-th and j-th group. $pdiff(g_{i})$ denotes the difference between the package list in the i-th group and all other groups.
@@
GValue Fuse. Based on the four dimensions score, we calculate the average value (i.e., $GValue(g_{i})$) as the comprehensive evaluation metric for assessing the quality of a group.
@@ VI-B1 Metrics Validation
We validated our GValue metric by evaluating five groups randomly selected from five different Linux distributions. The ground truth for these groups was established through a survey completed by 16 skilled participants. Then we calculate the Spearman correlation [37] between the human score and GValue score as an evaluation metric. The result is shown in Figure 9 and indicates the GValue have a strong correlation with human score ($rs>0.7$) in each aspect.
@@ VI-B2 Problems Analysis
Summary for RQ3: We develope a multi-dimensional feature-based method to evaluate group quality, which shows a strong correlation with human assessments. Besides, our analysis reveals that 16% of the groups exhibit poor quality, with two types of issues in the adoption of P2G.
@@ VII-A Methodology
VII-A1 Topic Analysis on group descriptions
Latent Dirichlet Allocation [38] (LDA) is a widely used unsupervised learning method for extracting topics from large-scale unstructured text. By mining the latent semantic structure of the text, LDA enables topic modeling across a collection of descriptions to identify common themes. Thus, we use LDA to conduct topic analysis on the group descriptions. Specifically, we start with preprocessing for the group descriptions, including tokenizing and removing stopwords using NLTK [39]. We then construct a vocabulary dictionary using the Gensim [40] and generate a group-word frequency matrix, representing the frequency of each word in each group. Finally, each topic is characterized by the top ten words with the highest weights, revealing the latent semantic structure in the group descriptions. Additionally, as LDA requires specifying the number of topics in advance, we perform a series of experiments with various numbers of topics and evaluate them using the coherence score to determine the optimal number.
@@ VII-A2 Extraction of Key Information from package descriptions
We extract and compare the key information from the descriptions of the packages in groups with those that are not. Package descriptions often contain many redundant and trivial words, so we perform tokenization and stopword removal. Then, we use the term frequency-inverse document frequency (TF-IDF) algorithm [41] to mine keywords. The TF-IDF algorithm determines the importance of a word by comparing its frequency in a specific package description against its rarity across all package descriptions. The algorithm can measure how relevant a word is within a given package description. Specifically, for a set of package descriptions $S$, a specific package description $p$ from this set, and a word $w$ in $p$, the relevance $i_{w}$ of the word is calculated as follows:
@@
where $f_{w,p}$ represents the frequency of $w$ in $p$, $\|S\|$ is the number of package descriptions in the set $S$, and $f_{w,S}$ denotes the frequency of $w$ across the entire set $S$. Using this algorithm, we extract the most relevant words from each package description. To help developers determine whether to apply P2G to a package, we collect and compare the keywords from package descriptions within and outside of groups.
@@ VII-B1 Topic in group descriptions
we set the number of topics ranging from 1 to 20 and calculate their coherence scores. We find the coherence score is the highest (i.e., 0.48) when the number is 3. Figure 10 presents the topics for group descriptions under this situation. We can observe that they mainly involve the following categories: Applications (e.g., fonts and tools), Development (e.g., programming and libraries), Graphical Desktop (e.g., desktop and KDE), Networks (e.g., network and virtualization), and Servers (e.g., servers). This suggests that developers tend to adopt P2G when a package’s functionality belongs to one of these categories.